The area of the rectangle shown is 63 in^2. Find the length and width of the rectangle.
Below this there is a rectangle with the length as (x+8) and the width as (2x+5)

Answer by rothauserc(4718)

Area(A) = length times width
A = (x+8) times (2x+5)
63 = 2x^2 + 21x + 40
2x^2 + 21x - 23 = 0
this factors into
(2x+23) times (x-1) = 0
x = -23/2 or 1
x = 1 is the correct answer
length = x + 8 = 9 inches
width = 2x + 5 = 7 inches
note that Area = 9 times 7 = 63 cm^2